# FedAGHN: Personalized Federated Learning with Attentive Graph HyperNetworks

### v1

@@ Abstract
Personalized Federated Learning (PFL) aims to address the statistical heterogeneity of data across clients by learning the personalized model for each client. Among various PFL approaches, the personalized aggregation-based approach conducts parameter aggregation in the server-side aggregation phase to generate personalized models, and focuses on learning appropriate collaborative relationships among clients for aggregation. However, the collaborative relationships vary in different scenarios and even at different stages of the FL process. To this end, we propose Personalized Federated Learning with Attentive Graph HyperNetworks (FedAGHN), which employs Attentive Graph HyperNetworks (AGHNs) to dynamically capture fine-grained collaborative relationships and generate client-specific personalized initial models. Specifically, AGHNs empower graphs to explicitly model the client-specific collaborative relationships, construct collaboration graphs, and introduce tunable attentive mechanism to derive the collaboration weights, so that the personalized initial models can be obtained by aggregating parameters over the collaboration graphs. Extensive experiments can demonstrate the superiority of FedAGHN. Moreover, a series of visualizations are presented to explore the effectiveness of collaboration graphs learned by FedAGHN.
@@ 1 Introduction
Federated Learning (FL) has provided a new paradigm for training the global model through client-side collaboration in distributed systems without sharing clients’ private data. However, the data distributions among different clients exhibit statistical heterogeneity in real-world scenarios, posing challenges for traditional FL approaches which rely on a single global model to adapt for the diverse local data of different clients Mills et al. (2021).
Personalized Federated Learning (PFL) offers a novel perspective to address statistical heterogeneity. The main concept is to learn a personalized model for each client rather than sharing a single global model. Among various PFL approaches, the personalized aggregation-based approach naturally fits the need of personalized adaptation for local data in different clients. In this work, we further investigate this approach with the focus on the collaboration mechanism.
The personalized aggregation-based approach performs personalized aggregation in the server-side aggregation phase instead of the global average aggregation. It mainly aims to obtain appropriate collaborative relationships for parameter aggregation, so that each client can receive additional gains without sharing local data. To achieve the goal, three key points may need to be considered:
@@
    Collaboration weights and weight differences. The collaboration weights offer a pattern for aggregating clients’ parameters and producing the personalized aggregated model. The weight differences of different FL scenarios should be adapted to the specific FL scenario considering its data distributions or other characteristics over clients.
@@
    Layer-wise collaboration differences. For deep learning models with multiple layers, personalized aggregation should consider the layer-wise personalization. Previous works LeCun et al. (2015); Ma et al. (2022); Lee et al. (2023) have suggested that different layers have different behaviours, e.g., shallow layers capture generic information, while deep layers acquire specific information.
Regarding the three points, most previous methods focus on certain aspect(s). For example, as shown in Figure 1, CFL Sattler et al. (2020) only considers collaboration with clustered intra-group collaborating clients; H-FedAMP Huang et al. (2021) considers the differences among collaborating clients, but the collaboration weight of the central client remains fixed. pFedLA Ma et al. (2022) and FedALA Zhang et al. (2023a) optimize trainable parameters to implicitly capture relationships, but encounter challenges in effectively capturing appropriate collaborative relationships.
Specifically, we propose a novel personalized aggregation-based FL method, Personalized Federated Learning with Attentive Graph HyperNetworks (FedAGHN), which employs the proposed Attentive Graph HyperNetworks (AGHNs) to generate client-specific personalized initial models. At each communication round, AGHNs maintain a set of collaboration graphs for each client and each layer. The collaboration graph is initialized by taking both personalized local model parameters and its updates as node features. The tunable attentive mechanism is introduced to calculate collaboration weights. It leverages prior knowledge from the updates of models and two trainable parameters optimized towards the client’s objective. One trainable parameter is employed to explicitly tune the weight differences of collaborating clients (Point 1), and another one is used to adjust the self-participation strength in collaboration (Point 2). After that, the personalized initial model of a client at each layer (Point 3) can be obtained by weighted aggregation over the corresponding learned graph.
@@
  • We introduce two sets of trainable parameters in FedAGHN, which is beneficial to capture the dynamic evolution of collaborative relationships. These trainable parameters are optimized dynamically in the FL process to align with the objective of each client for learning the client-specific personalized model.
  • We conduct extensive experiments and discover several meaningful insights by visualizing and analyzing the collaboration graphs from different points of view, such as overall patterns, layer-wise patterns, and the evolution of patterns during FL process.
@@ 2.1 Personalized Federated Learning
Statistical heterogeneity has become a crucial challenge in federated learning settings Meng et al. (2024); Tang et al. (2024). Personalized federated learning aims to learn models that better fit each client’s local data, enabling them to effectively tackle the statistical heterogeneity in FL Dinh et al. (2022); Li et al. (2021a). Recently, the development of PFL has led to the emergence of various approaches Tan et al. (2023); Zhang et al. (2023c), such as local adaptation, meta-learning-based Jiang et al. (2019); Fallah et al. (2020), multi-task-learning-based Marfoq et al. (2021), regularization-based T Dinh et al. (2020); Li et al. (2021b), parameter decoupling Collins et al. (2021); Chen and Chao (2022); Zhang et al. (2023b), knowledge distillation-based Wu et al. (2022); Xu et al. (2023) and personalized aggregation-based approaches Li et al. (2021c); Zhang et al. (2023a). Details of various PFL approaches are provided in Appendix A.
In this work, we focus on personalized aggregation-based methods. Specifically, CFL classifies clients into groups, each group sharing the same personalized model aggregated within the group. FedAMP and H-FedAMP address personalized aggregation with a client-wise approach using attention-based message passing mechanisms. pFedLA produces layer-wise aggregation weights and implicitly optimizes them via hypernetworks. FedALA learns collaboration matrices between local models and the global model for aggregation.
@@ 2.2 Graph Learning
Graph learning has attracted much attention in recent years Wang et al. (2022). The attention mechanism is introduced over the relational weights to be optimized with certain objectives Veličković et al. (2018); Zhang et al. (2024), such that the optimized weights better reflect the relations between nodes in terms of the objectives. In PFL, each client could have different specific relations to other clients due to inherent different data distributions, while the relations between clients are often unknown and complicated. It thus becomes critical to learn the relations between clients.
@@ 2.3 Hypernetworks
Hypernetwork is a technique that utilizes neural network to generate parameters (such as weights) for the target network, with the objective of enhancing its performance in target tasks. The property of hypernetworks, enabling the generation of improved client model parameters in the target task, makes them highly suitable for PFL Tang and Yao (2024). pFedHN Shamsian et al. (2021) maintains a unique embedding vector for each client, serving as input to the hypernetwork for generating personalized model parameters. pFedLA Ma et al. (2022) utilizes hypernetworks to generate layer-wise weighted aggregation parameters.
[…]
@@ 3 The Proposed Method
In this section, we first formally formulate the problem of PFL. After that, we present the overall framework and algorithm of the proposed FedAGHN. The technical details of AGHNs are finally elaborated step by step.
@@ 3.1 Problem Formulation
The objective of PFL is to learn a personalized local model for each client without sharing their private data. Assume that there are $N$ clients, and $\Theta=\left\{\theta_{1},\dots,\theta_{N}\right\}$ represents the parameters of these personalized local models, trained on their local datasets $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{N}\}$. The datasets $\mathcal{D}$ are sampled from Non-IID settings. The objective function of PFL can be formulated as
@@
The personalized aggregation-based methods mainly focus on the aggregation step to aggregate personalized models for each client, i.e.,
@@
where $\bar{\theta}_{i}^{(t+1)}$ denotes the personalized initial model of client $i$ in the next round $t+1$ and $\text{Agg}_{i}$ represents the aggregation function of different personalized aggregation-based methods, which is specified as the $\text{AGHN}_{i}$ in our method shown in Eq. (3).
@@ 3.2 FedAGHN Algorithm
In view of the fact that capturing the better fine-grained collaborative relationships among clients is the most essential and bottom-level requirement in the heterogeneous scenario of PFL, we propose a novel aggregation-based PFL method, namely Personalized Federated Learning with Attentive Graph HyperNetworks (FedAGHN), which utilizes our proposed Attentive Graph HyperNetworks (AGHNs) to generate client-specific personalized initial models.
The proposed FedAGHN aims to provide the collaboration graphs (i.e., collaboration weights) for the model aggregation among $N$ clients. FedAGHN also integrates prior knowledge to explicitly model the collaboration among clients, and it can adaptively adjust collaborative relationships through the optimization of hypernetworks’ parameters with respect to the clients’ objectives in an end-to-end fashion.
Specifically, we maintain a dedicated AGHN at the server side for each client, which means that there are $\text{AGHNs}=\{\text{AGHN}_{1},\text{AGHN}_{2},...,\text{AGHN}_{N}\}$ on the server. AGHNs are employed to capture the collaboration among clients, generating collaboration graphs $G=\{G_{1},G_{2},...,G_{N}\}$ for each client. Each graph $G_{i}$ also contains multiple subgraphs at the layer-wise level. Based on these layer-wise collaboration graphs, AGHNs then produce personalized initial models $\bar{\Theta}=[\bar{\theta}_{1},\bar{\theta}_{2},...,\bar{\theta}_{N}]$ which are dedicated to each client, rather than aggregating a single global model for all clients. In the server aggregation phase of each round $t$, the personalized initial model $\bar{\theta}^{(t+1)}_{i}$ in the next round $t+1$ for client $i$ can be obtained by
@@
where the input of $\text{AGHN}_{i}$ has two parts: the collection of personalized local model parameters denoted as ${{\Theta}^{(t)}}$, and the set of the updates of personalized local model parameters returned by the clients denoted as ${\Delta\Theta}^{(t)}$. Note that, $p_{i}$ and $q_{i}$ are the parameters of $\text{AGHN}_{i}$, which are vital for AGHNs to explicitly model and tune the collaborative relationships.
The overall framework of FedAGHN is illustrated in Figure 2, and its pseudocode is provided in Appendix B.
3.3 Attentive Graph Hyper-Networks
In this section, we take the attentive graph hypernetwork for client $i$ (i.e., $\text{AGHN}_{i}$) as an example to elaborate the technical details. For simplicity, we now focus on one communication round and remove the superscript reflecting the round. The illustration is shown in Figure 3.
The key of AGHNs is to learn the collaboration graphs (i.e., collaboration weights) such that each client can aggregate its beneficial model parameters from all clients based on the learned collaboration graphs. As the layers of deep learning models may have different functionalities, we also consider different collaboration patterns for different layers respectively, so that $\text{AGHN}_{i}$ for client $i$ maintains $R$ collaboration graphs $G_{i}=\{G^{1}_{i},G^{2}_{i},...,G^{R}_{i}\}$ where $R$ is the number of layers in the model at client $i$. Note that, there is no collaboration among different layers.
For convenience, we follow the graph learning terminology to present the attentive graph hyper-network at client $i$ and layer $r$, which corresponds to the collaboration graph $G_{i}^{r}$.
3.3.1 Initialize collaboration graphs
As for the initialization of graph topology, the collaboration graph $G^{r}_{i}$ is established for the central node (client) $i$, such that there are totally $N$ nodes with edges connected to node $i$ from all nodes $j$ (including self-connection from $i$).
Regarding the initialization of node features $(h^{r};a^{r})$, we take the personalized local model parameters $\Theta^{r}$ from clients at layer $r$ to set the node features $h^{r}$ as $h^{r}=\{h_{1}^{r},h_{2}^{r},...,h_{N}^{r}\}=\{\theta_{1}^{r},\theta_{2}^{r},....,\theta_{N}^{r}\}$ and we take the updates of personalized local model parameters $\Delta\Theta^{r}$ from clients at layer $r$ to set the node features $a^{r}$ as $a^{r}=\{a_{1}^{r},a_{2}^{r},...,a_{N}^{r}\}=\{\Delta\theta_{1}^{r},\Delta\theta_{2}^{r},....,\Delta\theta_{N}^{r}\}$ where the node features $h^{r}$ is employed for aggregation over graphs, and the node features $a^{r}$ is utilized to calculate attentive weights. The initial node features at the first round is randomly parameterized, whereas the initial node features at other rounds come from their previous round respectively.
3.3.2 Calculate attentive weights
Specifically, we calculate the collaboration weights between node $i$ and each node $j\in\{1,2,...,N\}$ via a newly proposed tunable attentive mechanism:
In order to optimize the attentive weights with respect to the client $i$’s objective, we introduce two trainable parameters to further tune the attentive weights. Concretely, $q^{r}_{i}$ is adopted to tune the attentive weights between central node $i$ and other nodes, while $p^{r}_{i}$ is used to tune the attentive weight of self-connection of central node $i$. Note that, the two trainable parameters $q^{r}_{i}$ and $p^{r}_{i}$ are vital for $\text{AGHN}_{i}$, since they can explicitly capture and tune the collaborative relationships according to the client’s objective in the end-to-end paradigm.
To accelerate convergence and avoid exploding gradients, the normalized attentive weights $\alpha_{i}^{r}=\{\alpha_{i1}^{r},\alpha_{i2}^{r},...,\alpha_{iN}^{r}\}$ for collaboration graph $G_{i}^{r}$ can be calculated via
$\alpha_{ij}^{r}=\frac{\widetilde{\alpha}_{ij}^{r}}{{\textstyle\sum_{j=1}^{n}}\widetilde{\alpha}_{ij}^{r}}\quad\forall j\in\{1,2,...,N\}.$ (5)
3.3.3 Aggregate parameters over graphs
The collaboration graph $G_{i}^{r}$ includes the attentive weights $\{\alpha_{i1}^{r},\alpha_{i2}^{r},...,\alpha_{iN}^{r}\}$ and the node features $h^{r}$ from previous round. The personalized initial model $\bar{\theta}_{i}^{r}$ for client $i$ at layer $r$ can be obtained by weighted aggregating over $G_{i}^{r}$ as follows:
${\bar{\theta}}_{i}^{r}=\sum_{j\in\{1,2,...,N\}}\alpha_{ij}^{r}{h_{j}^{r}}.$ (6)
3.3.4 The update of AGHN
Following the previous works Shamsian et al. (2021); Ma et al. (2022), we accordingly adopt a more general way to update $p_{i}$ and $q_{i}$ based on gradient descent, the details can be found in Appendix B.
4 Experimental Settings
This section briefly describes the experimental settings. More detailed experimental settings are presented in Appendix C.

### v2

@@ Abstract
Personalized Federated Learning (PFL) aims to address the statistical heterogeneity of data across clients by learning the personalized model for each client. Among various PFL approaches, the personalized aggregation-based approach conducts parameter aggregation in the server-side aggregation phase to generate personalized models, and focuses on learning appropriate collaborative relationships among clients for aggregation. However, the collaborative relationships vary in different scenarios and even at different stages of the FL process. To this end, we propose Personalized Federated Learning with Attentive Graph HyperNetworks (FedAGHN), which employs Attentive Graph HyperNetworks (AGHNs) to dynamically capture fine-grained collaborative relationships and generate client-specific personalized initial models. Specifically, AGHNs empower graphs to explicitly model the client-specific collaborative relationships, construct collaboration graphs, and introduce tunable attentive mechanism to derive the collaboration weights, so that the personalized initial models can be obtained by aggregating parameters over the collaboration graphs. Extensive experiments can demonstrate the superiority of FedAGHN. Moreover, a series of visualizations are presented to explore the effectiveness of learned collaboration graphs.
footnotetext: This is the authors’ final accepted manuscript of the article published in Knowledge-Based Systems, 2025. The published version is available at: https://doi.org/10.1016/j.knosys.2025.114355footnotetext: © 2025. This manuscript version is made available under the $C$ BY-NC-ND 4.0 license https://creativecommons.org/licenses/by-nc-nd/4.0/
@@ 1 Introduction
Data privacy and security concerns in the training of deep learning models, along with the phenomenon of isolated data islands, have attracted significant attention from researchers, government agencies, and other stakeholders. Federated Learning (FL) has provided a new paradigm for training the global model through client-side collaboration in distributed systems without the sharing of clients’ private data. The idea behind it is to transmit model parameters rather than private local data. For instance, FedAvg McMahan et al. (2017) allows clients to perform local training and then upload the update of local model parameters to the server for average aggregation, resulting in a global model that is subsequently distributed back to the clients. However, in real-world scenarios, the data distributions among different clients exhibit statistical heterogeneity, posing significant challenges for traditional FL approaches, which rely on a single global model, in adapting to the diverse local data of different clients Mills et al. (2021).
Personalized Federated Learning (PFL) offers a novel perspective to address statistical heterogeneity. The main concept is for each client to learn a personalized model, rather than sharing a single global model. It enables each client to learn a personalized local model by acquiring information from other clients and/or the global model. There exist various approaches to PFL, such as local adaptation, meta-learning-based Jiang et al. (2019); Fallah et al. (2020), multi-task-learning-based Marfoq et al. (2021), regularization-based T Dinh et al. (2020); Li et al. (2021b), parameter decoupling Collins et al. (2021); Chen and Chao (2022); Zhang et al. (2023b), knowledge distillation-based Wu et al. (2022); Xu et al. (2023) and personalized aggregation-based approaches Li et al. (2021c); Zhang et al. (2023a); He et al. (2024). Among these, the personalized aggregation-based approach naturally fits the need for personalized adaptation to the local data of different clients. Herein, client-specific personalized models are generated based on collaborative relationships among the clients. In this work, we further investigate this approach with a focus on the collaboration mechanism and collaborative relationship learning.
In the personalized aggregation-based approach, personalized aggregation is performed in the server-side aggregation phase instead of the global average aggregation. The aim is mainly to obtain appropriate collaborative relationships for parameter aggregation, such that each client can receive additional gains without sharing local data Ge et al. (2024). To achieve the goal, three key points may need to be considered:
@@
    Collaboration weights and weight differences. The collaboration weights offer a pattern for aggregating clients’ parameters and producing the personalized aggregated model. The weight differences of different FL scenarios should be adapted to the specific FL scenario considering its data distributions or other characteristics over the clients.
@@
    Layer-wise collaboration differences. For the deep learning models with multiple layers, personalized aggregation should consider the layer-wise personalization. Past studies LeCun et al. (2015); Ma et al. (2022); Lee et al. (2023) have suggested that different layers have different behaviours, e.g., the shallow layers capture generic information, whereas the deep layers acquire specific information.
With regard to the three aforementioned points, most existing methods focus only on certain aspect(s). For example, as shown in Figure 1, CFL Sattler et al. (2020) only considers collaboration with clustered intra-group collaborating clients; H-FedAMP Huang et al. (2021) considers the differences among collaborating clients, but the collaboration weight of the central client remains fixed. pFedLA Ma et al. (2022), FedALA Zhang et al. (2023a) and FedLGS He et al. (2024) optimize trainable parameters to implicitly capture relationships, but encounter challenges in effectively capturing appropriate collaborative relationships.
To fulfill all three key points, we propose a novel personalized aggregation-based FL method, Personalized Federated Learning with Attentive Graph HyperNetworks (FedAGHN), which employs the proposed Attentive Graph HyperNetworks (AGHNs) to generate client-specific personalized initial models. In each communication round, AGHNs maintain a set of collaboration graphs for each client and each layer. The collaboration graph is initialized by taking both personalized local model parameters and its updates as node features. A tunable attentive mechanism is introduced to calculate collaboration weights, leveraging prior knowledge from the updates of models and two trainable parameters optimized towards the client’s objective. One trainable parameter is employed to explicitly tune the weight differences of collaborating clients (Point 1), and another one is used to adjust the self-participation strength in the collaboration (Point 2). After that, the personalized initial model of a client at each layer (Point 3) can be obtained by weighted aggregation over the corresponding learned graph. These points are also the main innovations of our FedAGHN.
@@
  • We introduce two sets of trainable parameters in FedAGHN, which are beneficial for capturing the dynamic evolution of collaborative relationships. These trainable parameters are optimized dynamically in the FL process to align with the objective of each client for learning the client-specific personalized model.
  • We conduct extensive experiments and discover several meaningful insights by visualizing and analyzing the collaboration graphs from different points of view, such as overall patterns, layer-wise patterns, and the evolution of patterns during the FL process. The code is available at https://github.com/songjiarui00/FedAGHN2025.
@@ 2.1 Personalized Federated Learning
Statistical heterogeneity has become a crucial challenge in FL settings Ding et al. (2024); Zhang et al. (2025, 2024b).
PFL aims to learn models that better fit each client’s local data, enabling them to effectively tackle the statistical heterogeneity in FL Zheng et al. (2025); Li et al. (2021a). Recently, the development of PFL has led to the emergence of various approaches Tan et al. (2023); Zhang et al. (2023c). Vanilla PFL methods, such as FedAvg-FT and FedProx-FT, employ local training of the global model to obtain personalized models, which can be classified as local adaptation methods. On the other hand, meta-learning-based methods Jiang et al. (2019); Fallah et al. (2020) and multi-task-learning-based methods Marfoq et al. (2021) apply the principles of meta-learning and multi-task learning, respectivety, to PFL, aiming to enhance the model adaptation.
Regularization-based methods, such as pFedMe T Dinh et al. (2020) and Ditto Li et al. (2021b), have new ideas for using global model in PFL. These methods learn personalized models for each client and employ L2 regularization to constrain the distance between the personalized models and the global model. Another feasible way to absorb model information is by knowledge distillation. FedPAC Xu et al. (2023) is one of a number of knowledge distillation-based methods, which introduce global feature centroids for explicit local-global feature representation alignment.
Parameter decoupling methods are commonly employed to achieve better personalized or shared models. For example, FedRep Collins et al. (2021) first trains personalized classification layers and then trains a feature extractor, FedCP Zhang et al. (2023b) handles global and personalized information in the features separately using a global head and a personalized head, employing conditional policies to differentiate between global and personalized information. FedAS Yang et al. (2024) uses parameter alignment and client synchronization mechanisms to suppress inconsistencies between the personalized parameter part and the shared parameter part.
Personalized aggregation-based methods aim to obtain relationships among clients and leverage them as a basis for aggregation. Specifically, CFL Sattler et al. (2020) classifies clients into groups, with each group sharing the same personalized model aggregated within the group. FedAMP and H-FedAMP Huang et al. (2021) address personalized aggregation with a client-wise approach using attention-based message passing mechanisms. pFedLA Ma et al. (2022) produces layer-wise aggregation weights and implicitly optimizes them via hypernetworks, whereas FedALA Zhang et al. (2023a) learns collaboration matrices between local models and the global model for aggregation. FedLGS He et al. (2024) uses a multi-layer neural network to learn the aggregated weights of different clients.
Earlier-developed personalized aggregation-based methods focus mainly on certain aspects of collaboration among clients. By contrast, our work uses collaboration graphs to model the fine-grained collaborative relationships, and the collaboration weights on the graphs can then be adaptively optimized with clients’ objectives during the FL process.
Hypernetworks FedHN pFedLA FedAGHN
Number One HN shared by all clients One HN maintained per client One HN maintained per client
Personalization Client-wise Layer-wise Layer-wise
Architecture Neural network (large parameter size) Neural network (moderate parameter size) Lightweight graph neural network
Optimization Optimizing HN with client-specific embeddings Optimizing client-specific HN with embeddings Optimizing key tunable parameters in HN
Output Directly generating personalized model parameters Generating layer-wise aggregation coefficients for personalized model generation First constructing collaboration graphs; then generating personalized initial model parameters
Table 1: Comparison of Hypernetwork-based Methods
@@ 2.2 Graph Learning
Graph learning has attracted much attention in recent years Wang et al. (2022). The attention mechanism is introduced over the relational weights to be optimized with certain objectives Veličković et al. (2018); Zhang et al. (2024a), such that the optimized weights better reflect the relations between nodes in terms of the objectives. In PFL, each client could have different specific relations with other clients because of inherent different data distributions, and oftentimes, the relations between the clients are often unknown and complicated. It thus becomes critical to learn the relations between clients, so as to facilitate the model aggregation (i.e., collaboration) step and boost the performance of the local model.
The collaboration graph can be used to represent the relationship between clients in FL Zhang et al. (2022). However, natural graph structures do not exist in many real-world scenarios, thus requiring graph structure learning on the server side to establish collaboration graphs among clients Fu et al. (2022). A common approach is to construct collaboration graphs based on the model parameters of different clients. For example, SFL Chen et al. (2022) constructs a similarity matrix using the similarity of personalized model parameters, then it forms a graph with trainable parameters as edge weights, and optimizes the graph with regularization during local training at the client side; whereas pFedGraph Ye et al. (2023) designs optimization goals on the server side and proceeds to optimize the collaboration graph by solving a quadratic program. FedPnP Rasti-Meymandi et al. (2025) constructs graph relationships between clients based on label distribution and performs personalized collaboration according to the graph relationships.
Existing approaches thus far treat the model as a whole entity when constructing and utilizing the collaboration graph, disregarding the differences between different model layers. However, it is crucial to learn specific collaboration graphs for each model layer. By contrast, our work learns collaboration graphs for different model layers, these graphs can then be optimized via a tunable attentive mechanism during the FL process.
@@ 2.3 Hypernetworks
The hypernetwork technique utilizes neural network to generate parameters (such as weights) for the target network, with the objective of enhancing its performance in target tasks. This property of hypernetworks, enabling the generation of improved client model parameters in the target task, makes them highly suitable for PFL Tang and Yao (2024). For example, pFedHN Shamsian et al. (2021) maintains a unique embedding vector for each client, serving as input to the hypernetwork for generating personalized model parameters. By optimizing both the embedding vectors and hypernetwork parameters, pFedHN facilitates the generation of improved personalized model parameters. HAFL-GHN Litany et al. (2022) focuses on generating distinct model architecture weights for different clients. PLFL Wang (2025) proposes a lightweight hypernetwork to generate client model parameters and uses pruning strategies for personalized personalization. Meanwhile, rather than directly generating the complete weights of models, pFedLA Ma et al. (2022) utilizes hypernetworks to generate layer-wise weighted aggregation parameters. pFedLA has been verified to yield hypernetwork-generated aggregation parameters that can capture fine-grained relationships among clients, thereby enhancing the performance of personalized models. Additionally, the hypernetworks are on the server side, which mitigates the introduction of additional communication overhead.
The hypernetwork architecture in pFedLA employs SoftMax layers to generate the weights with differences for different model layers. On the other hand, in pFedHN, the majority of hypernetwork parameters are shared by all clients. The optimization process involves a larger number of parameters, which may potentially result in smaller differences between clients or between layers.
Regarding our work, we reduce the number of trainable parameters in the hypernetwork and retain only the vital training parameters that can tune the collaboration graphs, thus reducing the optimization burden while increasing the differences among model layers. A clearer comparison of the similarities and differences between our FedAGHN and these hypernetwork-based PFL methods is shown in the Table 1.
@@ 3 The Proposed Method
In this section, we first formally formulate the problem of PFL. After that, we present the overall framework and algorithm of the proposed FedAGHN. Finally, we elaborate on the technical details of AGHNs step by step.
@@ 3.1 Problem Formulation
The objective of PFL is to learn a personalized local model for each client without sharing their private data. If it is assumed that there are $N$ clients, and that $\Theta=\left\{\theta_{1},\dots,\theta_{N}\right\}$ represents the parameters of these personalized local models, trained on their local datasets $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{N}\}$, where the datasets $\mathcal{D}$ are sampled from Non-IID settings. The objective function of PFL can be formulated as
@@
Inherited from traditional federated learning, PFL also contains two main phases: local update and aggregation communication. For most PFL methods, the vanilla local update phase of the $t$-th round can be written as
$\theta_{i}^{(t)}=\theta_{i}^{(t)}-\eta\nabla_{\theta_{i}}\mathcal{L}_{i}(\theta_{i}^{(t)})$ (2)
where $\eta$ is the client learning rate and $\theta_{i}^{(t)}$ denotes the personalized local model of the $i$-th client in the $t$-th communication round.
Then in the aggregation phase, the server collects the updates of personalized local model parameters from each client and aggregates them to obtain a global model by
${\theta}_{\text{g}}^{(t+1)}=\sum_{i=1}^{N}\frac{|\mathcal{D}_{i}|}{|\mathcal{D}|}\theta_{i}^{(t)}$ (3)
where $\theta_{\text{g}}^{(t+1)}$ denotes the aggregated global model. This global model can be seen as having absorbed diverse information from heterogeneous clients in a sense.
After that, each client could use the aggregated global model as its initialization model in the next local update phase, i.e.,
$\bar{\theta}_{1}^{(t+1)}=\cdots=\bar{\theta}_{N}^{(t+1)}=\theta_{\text{g}}^{(t+1)}$ (4)
where $\bar{\theta}_{i}^{(t+1)}$ denotes the personalized initial model of client $i$ in the next round $t+1$ and set $\theta_{i}^{(t+1)}=\bar{\theta}_{i}^{(t+1)}$ in the beginning of the next local update phase.
However, due to the existence of data heterogeneity, the local models updated by Eq. (2) exhibit significant differences among various clients, which leads to the difficulty in obtaining a globally excellent model on all clients as shown in Eq. (3). Our AGHN method addresses this issue by adopting a personalized aggregation approach, designing a separate aggregation function for each client through graph hypernetworks.
Personalized aggregation-based methods focus mainly on the aggregation step to aggregate personalized models for each client, i.e.,
@@
where $\bar{\theta}_{i}^{(t+1)}$ denotes the personalized initial model of client $i$ in the next round $t+1$ and $\text{Agg}_{i}$ represents the aggregation function for different personalized aggregation-based methods. In our method, $\text{Agg}_{i}$ is specified as the $\text{AGHN}_{i}$, as shown in Eq. (6).
@@ 3.2 FedAGHN Algorithm
Given that capturing the better fine-grained collaborative relationships among clients is the most essential and bottom-level requirement in the heterogeneous scenario of PFL, we propose a novel aggregation-based PFL method, namely Personalized Federated Learning with Attentive Graph HyperNetworks (FedAGHN), which utilizes our proposed Attentive Graph HyperNetworks (AGHNs) to generate client-specific personalized initial models.
The proposed FedAGHN aims to provide the collaboration graphs (i.e., collaboration weights) for the model aggregation among $N$ clients. FedAGHN also integrates prior knowledge to explicitly model the collaboration among clients, and it can adaptively adjust collaborative relationships through the optimization of the hypernetworks’ parameters with respect to the clients’ objectives in an end-to-end fashion.
Specifically, we maintain a dedicated AGHN at the server side for each client, which signifies that there are $\text{AGHNs}=\{\text{AGHN}_{1},\text{AGHN}_{2},\dots,\text{AGHN}_{N}\}$ on the server. The AGHNs are employed to capture the collaboration among clients, generating collaboration graphs $G=\{G_{1},G_{2},\dots,G_{N}\}$ for each client. Each graph $G_{i}$ also contains multiple subgraphs at the layer-wise level. Based on these layer-wise collaboration graphs, the AGHNs then produce personalized initial models $\bar{\Theta}=[\bar{\theta}_{1},\bar{\theta}_{2},\dots,\bar{\theta}_{N}]$ which are dedicated to each client, rather than aggregating a single global model for all clients. In the server aggregation phase of each round $t$, the personalized initial model $\bar{\theta}^{(t+1)}_{i}$ in the next round $t+1$ for client $i$ can be obtained by
@@
where the input of $\text{AGHN}_{i}$ has two parts: the collection of personalized local model parameters denoted as ${\Theta}^{(t)}$}, and the set of the updates of personalized local model parameters returned by the clients denoted as ${\Delta\Theta}^{(t)}$. It should be noted that $p_{i}$ and $q_{i}$ are the parameters of $\text{AGHN}_{i}$, which are vital for AGHNs to explicitly model and tune the collaborative relationships.
The overall framework of FedAGHN is illustrated in Figure 2, and the pseudocode of FedAGHN algorithm is presented in Algorithm 1.